# What Matters in Data for DPO?

### v1

@@ 1 Introduction
The importance of aligning large language models (LLMs) with human preferences cannot be overstated. Two leading paradigms for achieving this alignment are Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022; Ouyang et al., 2022) and Direct Preference Optimization (DPO) (Rafailov et al., 2023). The key difference lies in whether an explicit reward model is trained (as in RLHF) or whether the model itself is optimized directly using preference data (as in DPO). Significant effort has been devoted to improving the performance of these methods by constructing more effective preference datasets. Common techniques include rejection sampling (i.e., generating multiple responses and selecting the best or worst, see Khaki et al. 2024), annotator rewriting/editing, and iterative use of on-policy data (Tajwar et al., 2024).
@@ 2 Preliminaries
Supervised Fine-tuning (SFT). SFT is is typically the first stage in adapting a pre-trained LLM to downstream tasks. Given a dataset $\mathcal{D}_{\text{SFT}}$ consisting of high-quality instruction-response pairs $(\mathbf{x},\mathbf{y})$ (Ouyang et al. 2022), the objective is to maximize the log-likelihood of the the demonstration data. Specifically, SFT is minimizing the loss function:
@@
Reinforcement Learning from Human Feedback (RLHF). After SFT, fine-tuning with human preference data is widely used to further align the model. RLHF begins by training a reward model $r(\mathbf{x},\mathbf{y})$ to reflect human preferences based on a preference dataset $\{(\mathbf{x}_{i},\mathbf{y}_{i.w},\mathbf{y}_{i,l})_{i\geq 1}\}$, where $\mathbf{y}_{i,w}$ and $\mathbf{y}_{i,l}$ denote the preferred and rejected response, respectively. The policy $\pi_{\theta}$ is then optimized, typically using reinforcement learning algorithms such as PPO (Schulman et al. 2017,Ouyang et al. 2022), by minimizing the following objective:
@@
Directed Preference Optimization (DPO). To simplify the process of RLHF, particularly to get rid of the training of a reward model, Rafailov et al. (2023) realize that the reward function can be represented by the learning policy:
@@
where $Z(\mathbf{x})=\sum_{\mathbf{y}}\pi_{\text{ref}}(\mathbf{y}|\mathbf{x})\exp{(r_{\bm{\theta}}(\mathbf{x},\mathbf{y})/\beta)}$ is the partition function. Based on the Bradley-Terry (BT) preference assumption (Bradley and Terry 1952), together with the pairs of chosen and rejected responses, DPO fine-tunes the language model by optimizing the following loss function:
@@
where $r^{*}$ is the true reward model. Rafailov et al. (2023) also derive the derivative of the DPO loss (4) with respect to the parameters $\bm{\theta}$:
@@ 3 Related Works
RL-based LLM Alignment and DPO. Following SFT, RL policy gradient methods are then employed to align model outputs with human preferences encoded through reward modeling. While early RL approaches like TRPO (Schulman et al., 2015) and PPO (Schulman et al., 2017) established foundational frameworks, their computational intensity motivated the development of resource-efficient alternatives such as RAFT (Dong et al., 2023), RRHF (Yuan et al., 2023), SLiC (Zhao et al., 2023), and ORPO (Hong et al., 2024). DPO (Rafailov et al., 2023) cleverly reinterprets the RL objective through contrastive loss by eliminating explicit reward value while maintaining stable training dynamics with reduced computational demands. The DPO framework has subsequently inspired multiple variants including KTO (Ethayarajh et al., 2024), IPO (Azar et al., 2024), and CPO (Xu et al., 2024a). Recent work by Shao et al. (2024) attempts to unify alignment-stage training paradigms through a generalized perspective.
@@
Data Quality in LLM Alignment. The critical role of data quality in LLM alignment has been rigorously established across training paradigms. Early studies (Zhou et al., 2023a) demonstrated its decisive impact in fine-tuning contexts, while the contemporary focus on reasoning (Muennighoff et al., 2025) further underscores the performance gains attainable through carefully curated alignment data. Empirical evidence specifically in DPO training (Morimura et al., 2024; Wu et al., 2024; Ivison et al., 2024) reveals two critical insights: (1) DPO exhibits stronger sensitivity to data quality compared to traditional RL methods like PPO; (2) strategic selection of high-quality samples improve DPO training performance. While Khaki et al. (2024) and Gou and Nguyen (2024) suggest larger preference gaps improve DPO, Pattnaik et al. (2024) and Xiao et al. (2025) find moderate gaps beneficial. However, a systematic understanding of the role of data quality remains lacking in the literature.
@@
On-policy DPO. Another fruitful stream of literature that is related to the proper usage of data is on-policy DPO implementations (Yuan et al., 2024; Chen et al., 2024; Guo et al., 2024; Rosset et al., 2024; Tajwar et al., 2024; Pang et al., 2024). Empirical analyses by Xu et al. (2024b) reveal that distributional mismatch between training data and the base model’s original domain disproportionately impacts DPO compared to PPO. On-policy DPO actively samples the intermediate model generations, which serves as an adaptive distributional bridge to mitigate out-of-domain degradation. Despite the potential benefit of on-policy DPO, excessive reliance on the on-policy data is very likely to induce training instability that can lead to a significant drop in model performance (Lambert et al., 2024; Deng et al., 2025). While trials have been made to balance the on-policy and off-policy data integration (Wang et al., 2025), understanding when and how on-policy data can be helpful also remains to be further explored.
@@ 4.1 The Role of Distributions of Chosen and Rejected Samples
  • DPO may deviate from RLHF. Rafailov et al. (2023) derive the DPO objective from RLHF by the observation the language model is secretly a reward model under the BT assumption. However, when the BT assumption does not necessarily hold, Eq. (8) implies that, in a more general sense, DPO implicitly performs reward learning with a specific reward function defined by $\pi_{w}$ and $\pi_{l}$:
@@
  • The quality of the chosen responses matters. Eq. (7) suggests that DPO performance is fundamentally limited by the quality of the chosen responses. More straightforwardly, when $\beta=1$ and the rejected samples are generated from the reference model, $\pi_{\text{DPO}}(\mathbf{y}|\mathbf{x})$ is just $\pi_{w}(\mathbf{y}|\mathbf{x})$. It is intuitively unreasonable to expect the language model generating the responses whose quality is much better than the chosen samples after DPO. Such an intuition has also been reflected in many DPO practices. For example, Dong et al. (2024b) showcase the effectiveness of using a response improver to polish what the current model generates and use it as the chosen samples.
  • The quality of the rejected responses may not always be critical. When $\pi_{l}(\mathbf{y}|\mathbf{x})$ and $\pi_{w}(\mathbf{y}|\mathbf{x})$ are very similar or even the same, DPO lacks a learning signal. However, the rejected distribution may not always take the fundamental role. As shown in Figure 1, imagine now we have two different distributions on the rejected samples $\pi^{\prime}_{l}(\mathbf{y}|\mathbf{x})$ and $\pi^{\prime\prime}_{l}(\mathbf{y}|\mathbf{x})$ that are only different from each other on the area where $\pi_{w}(\mathbf{y}|\mathbf{x})$ is small. Although $\pi^{\prime}_{l}(\mathbf{y}|\mathbf{x})$ and $\pi^{\prime\prime}_{l}(\mathbf{y}|\mathbf{x})$ are very different, the ratios ${\pi_{w}(\mathbf{y}|\mathbf{x})}/{\pi^{\prime}_{l}(\mathbf{y}|\mathbf{x})}$ and ${\pi_{w}(\mathbf{y}|\mathbf{x})}/{\pi^{\prime\prime}_{l}(\mathbf{y}|\mathbf{x})}$ can still be similar. Thus, $\pi^{\prime}_{l}(\mathbf{y}|\mathbf{x})$ and $\pi^{\prime\prime}_{l}(\mathbf{y}|\mathbf{x})$ may still lead to similar performances of DPO. In the literature, there have been some numerical results that implicitly implying such an idea. For example, Khaki et al. (2024) compare different preference data generation policies, including two based on $k$ generated answers. One is called Best-vs-worst where the chosen response is the best among the $k$ generated and the rejected response is the worst. The other one is Best-vs-random where the chosen response is again the best and the rejected response is randomly chosen from the rest k1k-1 responses. Interestingly, Khaki et al. (2024) report similar performances of Best-vs-worst and Best-vs-random across several different tasks. More numerical evidence is presented in our Section 5.
@@
  • The role of contrastiveness between chosen and rejected samples in DPO. Conventional wisdom in the field suggests that a larger preference gap between chosen and rejected responses enhances DPO training performance (Khaki et al., 2024; Gou and Nguyen, 2024). From Eq. (7), when $\pi_{w}(\mathbf{y}|\mathbf{x})$ and $\pi_{w}(\mathbf{y}|\mathbf{x})$ are nearly identical, DPO receives little useful signal to learn from. In this sense, greater contrastiveness helps avoid such degenerate cases. However, following what discussed above, once sufficient contrastiveness is achieved, further degrading the quality of the rejected responses may yield limited returns. A more fundamental benefit of contrastiveness appears to be that it encourages higher-quality chosen responses. This also interprets why increasing the number of candidates in rejection sampling tends to improve the performance: it increases the likelihood of selecting better chosen samples, while the specific choice of rejected responses is often less important—consistent with the findings in Khaki et al. (2024).
@@ 4.2 Online DPO with Fixed Chosen Samples
In current practice, the parameter $\beta$ is chosen between 0.03 and 0.1 under many circumstances, as seen in prior work such as Rafailov et al. (2023), as well as implementations from Cerebras.AI (Vishnevskiy, 2023) and Anyscale (Wang et al., 2024). In fact, through our dataset construction, it is also likely for $\epsilon_{\beta}$ to be positive. Therefore, even when $\epsilon_{\beta}$ is not negligibly small, the insights below can still be valid under many cases. The proof of Theorem 4.5 is based on the Taylor expansion of $\sigma(x)$, along with the fact that $\mathbf{y}_{w}$ and $\mathbf{y}_{l}$ are independently generated in $\bar{\mathcal{D}}_{\text{DPO}}$.
@@ 5.1 Experiment Settings
Base Model. In the following numerical experiments, we are utilizing Allen-AI’s open-sourced Llama-3.1-Tulu-3-8B-SFT checkpoint (Lambert et al., 2024) as the base model for DPO training (Sun et al., 2024). This model is exclusively supervised-finetuned (SFT) on a mix of publicly available and transparent SFT data from Meta’s official pre-trained model (Dubey et al., 2024), making it possible to guarantee no data overlap during the SFT and DPO stages.
Datasets. According to the data recipe of our base model, we select two public datasets, LAION-AI’s Open Assistant 2 (Köpf et al., 2023) and OpenBMB’s UltraFeedback (Cui et al., 2023), as the prompt datasets for our DPO training. We carefully curate the datasets to make sure the prompts are unique and not seen during the SFT stage.
Data Processing. For our experiments, we include multiple responses per prompt. For Open Assistant 2, we retain only first-turn dialogues and filter out prompts with fewer than 3 responses. For UltraFeedback, we consider two variants: the original version (ultrafeedback-original) (Cui et al., 2023), which provides 4 responses per prompt, and the tulu3 version (ultrafeedback-tulu3) (Lambert et al., 2024), which provides 2 responses per prompt. We focus on prompts that appear in both variants. After filtering, the Open Assistant 2 and UltraFeedback datasets contain 4,603 and 41,633 prompts, respectively. Besides the mentioned sources of responses, to ensure the abundance of the dataset for comparison, we also leverage the responses generated by the Mistral series model (Meng et al., 2024; Jiang et al., 2023). For each completion pair, i.e., a prompt and one of its responses, we use the Skywork-Reward-Gemma-2-27B-v0.2 (Liu et al., 2024) model as an oracle to assign quality scores. These scores serve as a proxy for data quality, enabling us to rank or categorize the samples and construct DPO datasets with different controlled qualities. We paired each prompt with five responses of varying quality, labeled as best, high, medium, low, and worst. We also use the filtered prompts of UltraFeedback dataset for on-policy response generation. For a detailed explanation, please refer to Appendix B.2.
Evaluation. To comprehensively assess the capabilities of our models, we employ a suite of standard evaluation benchmarks that measure diverse aspects of model performance. Based on established practices in the field, we include AlpacaEval-2 (Dubois et al., 2024), MMLU (Hendrycks et al., 2020), IFEval (Zhou et al., 2023b), TruthfulQA (Lin et al., 2021) and GSM8K (Cobbe et al., 2021) for estimating models’ abilities of general conversation, multitask understanding, instruction following, being truthful and informative, and mathematical reasoning, respectively.
@@ 5.3 Preference gap and exposure bias might not always be essential
Building on our understanding of the impact of response quality, we now turn to investigate two additional factors frequently discussed in the context of DPO training: preference gap and exposure bias. Conventional wisdom suggests that a larger preference gap between chosen and rejected responses enhances DPO training performance (Khaki et al., 2024; Gou and Nguyen, 2024) and that exposure bias arising from on-policy data also helps improve the model’s ability to learn preferences (Guo et al., 2024; Dong et al., 2024a). However, our findings respectively challenge these hypotheses, demonstrating that neither the preference gap nor exposure bias might not be as critical as previously believed. Instead, the quality of the chosen response emerges as the primary determinant of model performance, overshadowing the influence of these factors.
@@
Our investigation of exposure bias reveals a nuanced interaction between the inclusion of policy data and the quality of the chosen response. As shown in Table 4, introducing on-policy responses yields substantial gains in LC-AE2 and GSM8K metrics. However, such benefit strictly depends on the base data quality: low-quality configurations show minimal improvement despite equivalent on-policy proportions. This confirms that exposure bias mitigation only amplifies existing quality foundations rather than compensating for low-quality chosen responses. Notably, our implementation adopts the commonly used on-policy data ratios in the literature, as excessive reliance on such data is very likely to induce training instability that can lead to a significant drop in model performance (Lambert et al., 2024; Deng et al., 2025).

### v3

@@ 1 Introduction
The importance of aligning large language models (LLMs) with human preferences cannot be overstated. Two leading paradigms for achieving this alignment are Reinforcement Learning from Human Feedback (RLHF) (bai2022training; ouyang2022training) and Direct Preference Optimization (DPO) (rafailov2023direct). The key difference lies in whether an explicit reward model is trained (as in RLHF) or whether the model itself is optimized directly using preference data (as in DPO). Significant effort has been devoted to improving the performance of these methods by constructing more effective preference datasets. Common techniques include rejection sampling (i.e., generating multiple responses and selecting the best or worst, see khaki2024rs), annotator rewriting/editing, and iterative use of on-policy data (tajwar2024preference).
@@ 2 Preliminaries
Supervised Fine-tuning (SFT). SFT is is typically the first stage in adapting a pre-trained LLM to downstream tasks. Given a dataset $\mathcal{D}_{\text{SFT}}$ consisting of high-quality instruction-response pairs $(\mathbf{x},\mathbf{y})$ (ouyang2022training), the objective is to maximize the log-likelihood of the the demonstration data. Specifically, SFT is minimizing the loss function:
@@
Reinforcement Learning from Human Feedback (RLHF). After SFT, fine-tuning with human preference data is widely used to further align the model. RLHF begins by training a reward model $r(\mathbf{x},\mathbf{y})$ to reflect human preferences based on a preference dataset $\{(\mathbf{x}_{i},\mathbf{y}_{i.w},\mathbf{y}_{i,l})_{i\geq 1}\}$, where $\mathbf{y}_{i,w}$ and $\mathbf{y}_{i,l}$ denote the preferred and rejected response, respectively. The policy $\pi_{\theta}$ is then optimized, typically using reinforcement learning algorithms such as PPO (schulman2017proximal,ouyang2022training), by minimizing the following objective:
@@
Directed Preference Optimization (DPO). To simplify the process of RLHF, particularly to get rid of the training of a reward model, rafailov2023direct realize that the reward function can be represented by the learning policy:
@@
where $Z(\mathbf{x})=\sum_{\mathbf{y}}\pi_{\text{ref}}(\mathbf{y}|\mathbf{x})\exp{(r_{\bm{\theta}}(\mathbf{x},\mathbf{y})/\beta)}$ is the partition function. Based on the Bradley-Terry (BT) preference assumption (bradley1952rank), together with the pairs of chosen and rejected responses, DPO fine-tunes the language model by optimizing the following loss function:
@@
where $r^{*}$ is the true reward model. rafailov2023direct also derive the derivative of the DPO loss (4) with respect to the parameters $\bm{\theta}$:
@@ 3 Related Works
RL-based LLM Alignment and DPO. Following SFT, RL policy gradient methods are then employed to align model outputs with human preferences encoded through reward modeling. While early RL approaches like TRPO (schulman2015trust) and PPO (schulman2017proximal) established foundational frameworks, their computational intensity motivated the development of resource-efficient alternatives such as RAFT (dong2023raft), RRHF (yuan2023rrhf), SLiC (zhao2023slic), and ORPO (hong2024orpo). DPO (rafailov2023direct) cleverly reinterprets the RL objective through contrastive loss by eliminating explicit reward value while maintaining stable training dynamics with reduced computational demands. The DPO framework has subsequently inspired multiple variants including KTO (ethayarajh2024kto), IPO (azar2024general), and CPO (xu2024contrastive). Recent work by shao2024deepseekmath attempts to unify alignment-stage training paradigms through a generalized perspective.
@@
Data Quality in LLM Alignment. The critical role of data quality in LLM alignment has been rigorously established across training paradigms. Early studies (zhou2023lima) demonstrated its decisive impact in fine-tuning contexts, while the contemporary focus on reasoning (muennighoff2025s1) further underscores the performance gains attainable through carefully curated alignment data. Empirical evidence specifically in DPO training (morimura2024filtered; wu2024beta; ivison2024unpacking) reveals two critical insights: (1) DPO exhibits stronger sensitivity to data quality compared to traditional RL methods like PPO; (2) strategic selection of high-quality samples improve DPO training performance. While khaki2024rs and gou2024mixed suggest larger preference gaps improve DPO, pattnaik2024enhancing and xiao2025finding find moderate gaps beneficial. However, a systematic understanding of the role of data quality remains lacking in the literature.
@@
On-policy DPO. Another fruitful stream of literature that is related to the proper usage of data is on-policy DPO implementations (yuan2024selfreward; chen2024self; guo2024direct; rosset2024direct; tajwar2024preference; pang2024iterative). Empirical analyses by xu2024dpo reveal that distributional mismatch between training data and the base model’s original domain disproportionately impacts DPO compared to PPO. On-policy DPO actively samples the intermediate model generations, which serves as an adaptive distributional bridge to mitigate out-of-domain degradation. Despite the potential benefit of on-policy DPO, excessive reliance on the on-policy data is very likely to induce training instability that can lead to a significant drop in model performance (lambert2024t; deng2025less). feng2025pilaf propose PILAF, a theoretically-grounded sampling strategy for online and iterative DPO, which shares our work’s conceptual focus on attributing optimization signals to the data distribution. Their method uses policy interpolation to explicitly align the training gradient with the true oracle objective and help stablize the training process. While trials have also been made to balance the on-policy and off-policy data integration (wang2025inco), understanding when and how on-policy data can be helpful also remains to be further explored.
@@ 4.1 The Role of Distributions of Chosen and Rejected Samples
  • DPO may deviate from RLHF. rafailov2023direct derive the DPO objective from RLHF by the observation the language model is secretly a reward model under the BT assumption. However, when the BT assumption does not necessarily hold, Eq. (8) implies that, in a more general sense, DPO implicitly performs reward learning with a specific reward function defined by $\pi_{w}$ and $\pi_{l}$:
@@
  • The quality of the chosen responses matters. Eq. (7) suggests that DPO performance is fundamentally limited by the quality of the chosen responses. More straightforwardly, when $\beta=1$ and the rejected samples are generated from the reference model, $\pi_{\text{DPO}}(\mathbf{y}|\mathbf{x})$ is just $\pi_{w}(\mathbf{y}|\mathbf{x})$. It is intuitively unreasonable to expect the language model generating the responses whose quality is much better than the chosen samples after DPO. Such an intuition has also been reflected in many DPO practices. For example, dong2024self showcase the effectiveness of using a response improver to polish what the current model generates and use it as the chosen samples.
  • The quality of the rejected responses may not always be critical. When $\pi_{l}(\mathbf{y}|\mathbf{x})$ and $\pi_{w}(\mathbf{y}|\mathbf{x})$ are very similar or even the same, DPO lacks a learning signal. However, the rejected distribution may not always take the fundamental role. As shown in Figure 1, imagine now we have two different distributions on the rejected samples $\pi^{\prime}_{l}(\mathbf{y}|\mathbf{x})$ and $\pi^{\prime\prime}_{l}(\mathbf{y}|\mathbf{x})$ that are only different from each other on the area where $\pi_{w}(\mathbf{y}|\mathbf{x})$ is small. Although $\pi^{\prime}_{l}(\mathbf{y}|\mathbf{x})$ and $\pi^{\prime\prime}_{l}(\mathbf{y}|\mathbf{x})$ are very different, the ratios ${\pi_{w}(\mathbf{y}|\mathbf{x})}/{\pi^{\prime}_{l}(\mathbf{y}|\mathbf{x})}$ and ${\pi_{w}(\mathbf{y}|\mathbf{x})}/{\pi^{\prime\prime}_{l}(\mathbf{y}|\mathbf{x})}$ can still be similar. Thus, $\pi^{\prime}_{l}(\mathbf{y}|\mathbf{x})$ and $\pi^{\prime\prime}_{l}(\mathbf{y}|\mathbf{x})$ may still lead to similar performances of DPO. In the literature, there have been some numerical results that implicitly implying such an idea. For example, khaki2024rs compare different preference data generation policies, including two based on $k$ generated answers. One is called Best-vs-worst where the chosen response is the best among the $k$ generated and the rejected response is the worst. The other one is Best-vs-random where the chosen response is again the best and the rejected response is randomly chosen from the rest k1k-1 responses. Interestingly, khaki2024rs report similar performances of Best-vs-worst and Best-vs-random across several different tasks. More numerical evidence is presented in our Section 5.
@@
  • The role of contrastiveness between chosen and rejected samples in DPO. Conventional wisdom in the field suggests that a larger preference gap between chosen and rejected responses enhances DPO training performance (khaki2024rs; gou2024mixed). From Eq. (7), when $\pi_{w}(\mathbf{y}|\mathbf{x})$ and $\pi_{w}(\mathbf{y}|\mathbf{x})$ are nearly identical, DPO receives little useful signal to learn from. In this sense, greater contrastiveness helps avoid such degenerate cases. However, following what discussed above, once sufficient contrastiveness is achieved, further degrading the quality of the rejected responses may yield limited returns. A more fundamental benefit of contrastiveness appears to be that it encourages higher-quality chosen responses. This also interprets why increasing the number of candidates in rejection sampling tends to improve the performance: it increases the likelihood of selecting better chosen samples, while the specific choice of rejected responses is often less important—consistent with the findings in khaki2024rs.
@@ 4.2 Online DPO with Fixed Chosen Samples
In current practice, the parameter $\beta$ is chosen between 0.03 and 0.1 under many circumstances, as seen in prior work such as rafailov2023direct, as well as implementations from Cerebras.AI (vishnevskiy2023dpo) and Anyscale (wang2024dpo). In fact, through our dataset construction, it is also likely for $\epsilon_{\beta}$ to be positive. Therefore, even when $\epsilon_{\beta}$ is not negligibly small, the insights below can still be valid under many cases. The proof of Theorem 4.5 is based on the Taylor expansion of $\sigma(x)$, along with the fact that $\mathbf{y}_{w}$ and $\mathbf{y}_{l}$ are independently generated in $\bar{\mathcal{D}}_{\text{DPO}}$.
@@ 5.1 Experiment Settings
Base Model. In the following numerical experiments, we are utilizing Allen-AI’s open-sourced Llama-3.1-Tulu-3-8B-SFT checkpoint (lambert2024t) as the base model for DPO training (sun2024block). This model is exclusively supervised-finetuned (SFT) on a mix of publicly available and transparent SFT data from Meta’s official pre-trained model (dubey2024llama), making it possible to guarantee no data overlap during the SFT and DPO stages.
Datasets. According to the data recipe of our base model, we select two public datasets, LAION-AI’s Open Assistant 2 (kopf2023openassistant) and OpenBMB’s UltraFeedback (cui2023ultrafeedback), as the prompt datasets for our DPO training. We carefully curate the datasets to make sure the prompts are unique and not seen during the SFT stage.
Data Processing. For our experiments, we include multiple responses per prompt. For Open Assistant 2, we retain only first-turn dialogues and filter out prompts with fewer than 3 responses. For UltraFeedback, we consider two variants: the original version (ultrafeedback-original) (cui2023ultrafeedback), which provides 4 responses per prompt, and the tulu3 version (ultrafeedback-tulu3) (lambert2024t), which provides 2 responses per prompt. We focus on prompts that appear in both variants. After filtering, the Open Assistant 2 and UltraFeedback datasets contain 4,603 and 41,633 prompts, respectively. Besides the mentioned sources of responses, to ensure the abundance of the dataset for comparison, we also leverage the responses generated by the Mistral series model (meng2024simpo; jiang2023mistral7b). For each completion pair, i.e., a prompt and one of its responses, we use the Skywork-Reward-Gemma-2-27B-v0.2 (liu2024skywork) model as an oracle to assign quality scores. These scores serve as a proxy for data quality, enabling us to rank or categorize the samples and construct DPO datasets with different controlled qualities. We paired each prompt with five responses of varying quality, labeled as best, high, medium, low, and worst. We also use the filtered prompts of UltraFeedback dataset for on-policy response generation. For a detailed explanation, please refer to Appendix B.2.
Evaluation. To comprehensively assess the capabilities of our models, we employ a suite of standard evaluation benchmarks that measure diverse aspects of model performance. Based on established practices in the field, we include AlpacaEval-2 (dubois2024length), MMLU (hendrycks2020measuring), IFEval (zhou2023instruction), TruthfulQA (lin2021truthfulqa) and GSM8K (cobbe2021training) for estimating models’ abilities of general conversation, multitask understanding, instruction following, being truthful and informative, and mathematical reasoning, respectively.
@@ 5.3 Preference gap and exposure bias might not always be essential
Building on our understanding of the impact of response quality, we now turn to investigate two additional factors frequently discussed in the context of DPO training: preference gap and exposure bias. Conventional wisdom suggests that a larger preference gap between chosen and rejected responses enhances DPO training performance (khaki2024rs; gou2024mixed) and that exposure bias arising from on-policy data also helps improve the model’s ability to learn preferences (guo2024direct; dong2024rlhf). However, our findings respectively challenge these hypotheses, demonstrating that neither the preference gap nor exposure bias might not be as critical as previously believed. Instead, the quality of the chosen response emerges as the primary determinant of model performance, overshadowing the influence of these factors.
@@
Our investigation of exposure bias reveals a nuanced interaction between the inclusion of policy data and the quality of the chosen response. As shown in Table 4, introducing on-policy responses yields substantial gains in LC-AE2 and GSM8K metrics. However, such benefit strictly depends on the base data quality: low-quality configurations show minimal improvement despite equivalent on-policy proportions. This confirms that exposure bias mitigation only amplifies existing quality foundations rather than compensating for low-quality chosen responses. Notably, our implementation adopts the commonly used on-policy data ratios in the literature, as excessive reliance on such data is very likely to induce training instability that can lead to a significant drop in model performance (lambert2024t; deng2025less).